OpenStudy (anonymous):

sin(2arctan(4))=? I used to just make a triangle and evaluate it from there but not sure how to go about that with the "2". How should I solve this?

OpenStudy (anonymous):

WELCOME TO OPENSTUDY where u can ask questions and answer questions and give medals and fan people that helped you so much that want to ask a question you just have to do this @_Stephen_324 to get them to your question!!!!

OpenStudy (anonymous):

...

jimthompson5910 (jim_thompson5910):

Let \(\Large \theta = \arctan(4)\) Use an identity to get \[\Large \sin(2\arctan(4)) = \sin(2\theta)\] \[\Large \sin(2\arctan(4)) = 2\sin(\theta)\cos(\theta)\] \[\Large \sin(2\arctan(4)) = 2\sin(\arctan(4))\cos(\arctan(4))\]

jimthompson5910 (jim_thompson5910):

Hopefully you see what to do from here

OpenStudy (anonymous):

Thank You!

jimthompson5910 (jim_thompson5910):

you're welcome